OpenStudy (anonymous):

\[B=\left[\begin{matrix}3 & 3 \\ -1 & -1\end{matrix}\right]\] Find K if \[B^2\]=kB

OpenStudy (unklerhaukus):

k must be a matrix right?

OpenStudy (anonymous):

or a scaler

OpenStudy (unklerhaukus):

should we find B^2

OpenStudy (unklerhaukus):

\[\textbf B=\left[\begin{array}\ 3&3\\-1&-1 \end{array}\right]\] \[\textbf B^2=\left[\begin{array}\ 3&3\\-1&-1 \end{array}\right] \left[\begin{array}\ 3&3\\-1&-1 \end{array}\right]\]

OpenStudy (anonymous):

I think k is a scaler of 2

OpenStudy (anonymous):

Yep u r very correct @Hollywood_chrissy

OpenStudy (anonymous):

AS,